Find the solution using the elimination method
3x-y=8
x-1/3y=2

Answer by stanbon(75887)

Find the solution using the elimination method
3x-y=8
x-1/3y=2
----
Modify::
3x - y = 8
3x - y = 6
-------
That is a contradiction.
----
No solution
Cheers,
Stan H.